OpenStudy (anonymous):

There are many tangent lines to the graph of f(x) = (x-2)^1/2 Find equations of tangent lines that pass through the point (5,2). Note the point (5,2) is not on the graph of f. I need a hint. I know its going to be some (a,b) on the graph that has tangent line going through it s.t. it also passes thro. (5,2).

OpenStudy (asnaseer):

find the derivative first - this will give you the slope of the tangent line at any point of the graph. then find the value of the slope when x=5. you should hopefully be able to solve the rest yourself.

OpenStudy (anonymous):

Wai Wait.

OpenStudy (anonymous):

Ok this is my attempt. Please let me post it out.

OpenStudy (anonymous):

f(x) = (x-2)^{1/2}, then the derivative of f is: f'(x) = 1/2(x-2)^{-1/2}

OpenStudy (anonymous):

I have one major question

OpenStudy (anonymous):

When you have the tangent line equation, is it going to be: y = y'*(x) + b? I know that we have to sub in (5,2) first to get the y-int, then we have to sub into the equation (a,(a-2)^1/2), BUT I have a question, is it going to be: y = 1/2(a-2)^1/2*[a] + (whatever the y-int is) OR y = 1/2(a-2)^1/2 +(whatever the y-int is)

OpenStudy (anonymous):

Please let me know if you don't understand what I am trying to say.

OpenStudy (asnaseer):

sory - I just read the bit "Note the point (5,2) is not on the graph of f. " - so my suggestion above will not work. let me ready what you have written so far...

OpenStudy (anonymous):

I am saying that do you write in the form: y= y'*(x) + b OR y= y' + b <-- its because y' also has x-term in it, so I was confused).

OpenStudy (asnaseer):

equation of a straight line is: y = mx +c where m is the slope and c is the y-intercept

OpenStudy (asnaseer):

you have found the equation of the set of "slopes" by finding the derivative of the graph: \[f'(x) = 1/2(x-2)^{-1/2}\]

OpenStudy (asnaseer):

you also know that the tangent line must pass thru (5,2). so: 5 = m*2 + c giving c = 5-2m

OpenStudy (asnaseer):

so we can write the equation of the line as: y = mx + (5-2m) = m(x-2) + 5 this gives: m = (y-5)/(x-2) so I believe you should be able to equate this to the derivative above to get the set of lines asked for

OpenStudy (asnaseer):

i.e.\[m=f'(x)\]

OpenStudy (asnaseer):

am I making any sense to you or have I completely mis-understood the problem?

OpenStudy (asnaseer):

sorry - I got the x and y mixed up above: you also know that the tangent line must pass thru (5,2). so: 2 = m*5 + c giving c = 2-5m

OpenStudy (anonymous):

Ok let me tell you in words as to what my procedure is : 1) y = mx + b, where m = y' 2) use y = y'*x + b and sub in (5,2) to find y-int. 3) now we know there is a point (a,(a-2)^1/2) on curve s.t. its tangenet line crosses thro. (5,2) 4) Sub in (a,(a-2)^1/2 into the tangent line equation and solve for a.

OpenStudy (asnaseer):

so we can write the equation of the line as: y = mx + (2-5m) = m(x-5) + 2 this gives: m = (y-2)/(x-5)

OpenStudy (asnaseer):

yes - you are correct

OpenStudy (anonymous):

But it gets so ugly. How am I suppose to solve for a with all those radicals around?

OpenStudy (jamesj):

You know the equation of any tangent line: you have its slope and you know one point on the line; namely the point on the graph of f to which it is tangent. Write down the general equation of a tangent line. Then see whether or not it passes through (5,2)

OpenStudy (asnaseer):

I end up with:\[y=2+\frac{x-5}{2\sqrt{x-2}}\]

OpenStudy (jamesj):

f(x) = (x-2)^1/2 , Now for a point x = a, f'(a) = (1/2)(a-2)^(-1/2). The corresponding point on the graph is (a,f(a)) = (a,(a-2)^1/2). Hence the equation of the general tangent line is y - f(a) = f'(a) ( x - a ), i.e., \[ y - \sqrt{a-2} = \frac{1}{2\sqrt{a-2}} (x - a) \] Now this equation goes through (5,2) if \[ 2 - \sqrt{a-2} = \frac{1}{2\sqrt{a-2}} (5 - a) \] Multiply both sides by \( 2\sqrt{a-2} \), \[ 4\sqrt{a-2} - 2(a-2) = 5 - a \] i.e. \[ 4 \sqrt{a-2} = a + 1 \] Now square both sides and solve for a.

OpenStudy (anonymous):

JamesJ you are amazing. You just reminded me that I could have just used point-slope form.

OpenStudy (asnaseer):

@JamesJ - where did I go wrong?

OpenStudy (asnaseer):

was one of my assumptions incorrect?

OpenStudy (jamesj):

@asnaseer, you were right. You just didn't use the form for y = f(x) in your last equation.

OpenStudy (asnaseer):

aaahhhh - ok - now I see - thanks for the clarification @JamesJ